OpenStudy (anonymous):

A trapezoid has height h.One base is 2 units longer than the height. The other base is 3 times longer than the height . Write a polynomial for the area of the trapezoid. A.4h^2+2h B.2h^2+h. C.2h^2+1 D.2h^2+2h

OpenStudy (btaylor):

The area of a trapezoid is: \[A =\frac{1}{2}h(b_1+b_2)\] The problem gives us b1=h+2 and b2 = 3h. Therefore: \[A = \frac{1}{2}h(h+2+3h) = \frac{1}{2}h(4h+2) = \frac{1}{2}(4h^2+2h)=2h^2+h\] So your answer is B.

OpenStudy (anonymous):

Thanks

OpenStudy (btaylor):

yw